OpenStudy (anonymous):

If cos of theta = -4/9 with theta in quadrant 2, find sin of theta.

OpenStudy (jdoe0001):

what are the signs for cosine and sine in the 2nd quadrant anyway? both positive, both negative cosine positive and sine negative or viceversa?

OpenStudy (anonymous):

sines positive cosines negitive

OpenStudy (anonymous):

\[\sin \left(\cos ^{-1}\left(-\frac{4}{9}\right)\right)=\frac{\sqrt{65}}{9} \]

OpenStudy (jdoe0001):

\(\bf cos(\theta)=-\cfrac{4}{9}\to \cfrac{adjacent\ side}{hypotenuse}\to \cfrac{a=-4}{c=9}\)  \(\bf c^2=a^2-{\color{blue}{ b}}^2\implies \pm \sqrt{c^2-a^2}={\color{blue}{ b}} \\ \quad \\ \textit{we know that sine is positive, so we use the + of the root } \\ \quad \\ +\sqrt{c^2-a^2}={\color{blue}{ b}}\qquad recall\implies sin(\theta)=\cfrac{opposite\ side}{hypotenuse}\to \cfrac{{\color{blue}{ b}}}{c}\)

OpenStudy (anonymous):

so the answer is sqrt(97)/9?

OpenStudy (jdoe0001):

hmm

OpenStudy (anonymous):

or it is negative because sin is in quadrant 2 so -sqrt(97)/9 ???

OpenStudy (jdoe0001):

\(\bf c^2=a^2-{\color{blue}{ b}}^2\implies \pm \sqrt{c^2-a^2}={\color{blue}{ b}}\implies +\sqrt{(9)^2-(-4)^2}={\color{blue}{ b}} \\ \quad \\ \sqrt{81-16}={\color{blue}{ b}}\qquad sin(\theta)=\cfrac{\sqrt{81-16}}{9}\)

OpenStudy (jdoe0001):

"sines positive cosines negitive" <--- sine is postive, cosine is negative

OpenStudy (anonymous):

oooooh so sqrt(65)/9

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

ok thanks

OpenStudy (jdoe0001):

yw